Introduction to probabilistic seasonal forecasts

Michael K. Tippett

International Research Institute for Climate and Society The Earth Institute, Columbia University

ERFS Climate Predictability Tool Training Workshop May 4-9, 2009

Outline

▶ Physical basis. ▶ Ensembles. ▶ Empirical. ▶ MME.

Summary

▶
▶